OpenStudy (anonymous):

Simplify: 12x^4 ----- 18x^6 (the ----- means divided by)

OpenStudy (anonymous):

2/(3x^2)

OpenStudy (anonymous):

2/3X^2

OpenStudy (anonymous):

How do you get there though?

OpenStudy (anonymous):

divide both top and bottom parts by 3x^4

OpenStudy (anonymous):

12/18=2/3 X^4/X^6=1/X^2

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

first simplify 12/18 you get 2/3 then since its x^4/x^6 =x^4-6 x^-2 the answer is 3/2x^2